7) Derek Jeter currently has an investment portfolio that contains four stocks with a total value equal to $80,000. The portfolio has a beta (B) equal to 1.4. Derek wants an additional $20,000 in a stock that has B-2.4. After Derek ad portfolio, what will be the portfolios beta? to invest As the capital budgeting director for Fitzpatrick Coffins Company, you are evaluating construction of a new plant. The plant has a net cost of $5 million in Year O (today), and it will provide net cash inflows of S1 million at the end of Year 1, $1.5 million at the end of Year 2, and $2 million at the end of Years 3 throughS. If your required rate of return is 13%, what is the plants Payback, NPV and IRR? 8)
Answer #1

7 ) the four stocks with a total value equal = 80000

the portfolio = 1.4

the additional invest = 20000

the stock = 2.4

the portfolio beta = ( 80000 * 1.4 + 20000 * 2.4 ) / ( 80000 + 20000 )

= 160000 / 100000

= 1.6

8 )

CF = -5,1,1.5,2,2,2

Payback = 3 + 0.5 / 2 = 3.25 years

NPV = -5 + 1/(1+13%) + 1.5/(1+13%)2 + 2/(1+13%)3 + 2/(1+13%)4 + 2/(1+13%)5 = 0.7579

IRR = IRR({-5,1,1.5,2,2,2})

= 18.37%